Circular No. 8937

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA SAGITTARII 2008 S. Nakano, Sumoto, Japan, reports the discovery of a possible nova

(mag 8.4) on two 30-s unfiltered CCD survey frames taken on Apr. 18.784 UT by K. Nishiyama (Kurume, Fukuoka-ken, Japan) and F. Kabashima (Miyaki-cho, Saga-ken, Japan) using a 105-mm f/5.6 camera lens (limiting mag ∼ 12.6). An unfiltered CCD image taken with a 0.40-m f/9.8 reflector on Apr. 18.809 yields mag 8.4 and the following precise position for the new object: α = 18^{h}05^{m}58^{s}.88, δ = −27^{o}13′56′′.0 (equinox 2000.0). Nothing is visible at this position on the Digitized Sky Survey (DSS). A nearby USNO-B1.0 star has I mag 18.3 and position end figures 58^{s}.38, 55′′.9. Following posting on the Central Bureau’s unconfirmed-objects webpage, S. Dvorak, Clermont, FL, U.S.A., writes that his CCD exposures taken with a 0.25-m Meade LX200 reflector (+ ST-9XE camera) on Apr. 19.36 yield position end figures 58^{s}.85, 57′′.0 for the new variable. Dvorak adds that nothing appears at this position on a DSS red image from 1991, noting that nearby stars are obviously visible therein that have USNO-B1.0 red magnitudes around 20; he also remarks that a couple of stars at red mag ∼ 15 nicely bracket the apparent nova. Additionally, E. Guido and G. Sostero (Remanzacco, Italy) report that their confirming CCD images taken remotely with a 0.25-m f/3.4 reflector near Mayhill, NM, on Apr. 19.39 show the new object at position figures 58^{s}.90, 56′′.3; comparison with a U.K. Schmidt red plate obtained on 1996 Sept. 8 shows nothing at this position (limiting magnitude near 20). H. Yamaoka, Kyushu University, writes that K. Haseda (Toyohashi, Aichi, Japan) adds that M. Fujii (Kurashiki, Okayama, Japan) took a low-resolution spectrum with his 0.28-m reflector on Apr. 19.82, revealing Balmer lines and several broad absorption lines; the Hα line shows a prominent P-Cyg profile, which suggests that it is a genuine classical nova around or soon after maximum. The following additional CCD magnitudes have been reported for the nova (unfiltered unless otherwise noted): Apr. 11.593, [11 (P. Camilleri, Hurstville, Sydney, N.S.W.); 13.3, [11.0: (W. Liller, Viña del Mar, Chile; Tech Pan films with an orange filter, 85-mm camera lens); 13.765, [12.2 (Nishiyama and Kabashima); 14.805, [12.8 (Nishiyama and Kabashima); 15.743, [11.5 (K. Haseda, Toyohashi, Aichi, Japan, Canon EOS-5D digital camera + 120mm-f.l. lens; communicated by Yamaoka); 16.22, [11.0: (Liller); 19.36, V = 8.94, B = 9.98 (Dvorak; revision to CBET 1342); 19.39, B = 9.84, V = 8.90, R = 8.30 (Guido and Sostero; uncertainty ± 0.05 mag); 19.725, 7.9 (Nishiyama and Kabashima). Visual magnitude estimate by A. Amorim, Florianopolis, Brazil: Apr. 23.099, 6.5.

2008 April 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green